INAF queue observing, 2015 Oct 02-Oct 03

Observer: F. Cusano, R. Carini
Telescope Operator: Geno Bechetti
Support Astronomer: Olga Kuhn
Instrument: LBC, MODS

Summary

At the sunset the sky is cloud-covered, the wind speed is slighlty below the limit.
01:47 We stay closed.
03:23 sky is clear we open the dome, we start with LBC

Night duration (12\xB0 el twilight) 19:30-4:50 ( 02:30-11:50 UT) = 9 h 20m

Observing time: * 0 h*

Weather time loss: 5 hrs 10m (55%)

technical time loss 4 hrs 10m (45%)

Project_RUN Objects Filter ExpTime Notes
         

Details

03:40 start with LBC

03:45 we slew to Mignani B3
  • 03:46 blue primary mirror panic, we wait for the recovery
  • 03:53 sending again the preset, B3 copointing
  • dofpia, problem with the red channel, is not converging we apply manual correction to Z8,
  • 4:19 clouds coming in
  • 4:29 problem with red channel not solved and clouds coming in, we switch to MODS

04:30 switching to MODS

04:39 clouds cover all the sky

08:40 we open the dome, the sky is clear and humidity is decreasing

08:45 problem, the door of the left side is not opening
Topic revision: r6 - 05 Oct 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback